HUBBLE SPACE TELESCOPE DAILY REPORT #2042 PERIOD COVERED: 0000Z (UTC) 01/16/98 - 0000Z (UTC) 01/20/98 Daily Status Report as of 020/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/3 7815 (NICMOS Earth Flats) The Near Infrared Camera (Camera Three) was used to make a Earth flat field calibration image. This proposal obtains a series of earth flats for a subset of narrow and medium band NIC/3 filters in advance of the NIC/3 campaign. These flats will be used to generate the master flats for narrow-band filters and to determine the OTA illumination pattern as a function of wavelength. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Eight Sets of NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around stars GL-806, GM-Aur, HL-Tau, SVS-13, DS Tau GI-Tau, IP-Tau and DG-Tau to search for water in their disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The seventh iteration of this proposal was scheduled while NICMOS was in suspend. Then in the final iteration the acquisition default mentioned in 2.1 occurred. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.4 Completed Ten Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Eight Sets of NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of clusters GL-195A, GL-328, GL-213, GJ-1111, IC248, BD+010D4774, GL-83.1 and IC348. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The third and fourth iterations of this proposal was scheduled while NICMOS was in suspend. Otherwise, the observations were executed as scheduled, and no further anomalies were reported. 1.6 Completed Ten Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The eighth iteration of this proposal was scheduled while NICMOS was in suspend. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed NIC/1/2/3 7246 (Line Imaging of the L1448-Star Shock Region) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the star shock region of L1448-4. The bow shock associated with the infrared source L1448-4 star is one of the best examples of the interaction of an outflow from a recently formed star with the surrounding medium. As mentioned in 2.1, one of the acquisitions for the latter portion of this proposal succeeded on the third try. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed NIC/1/2/3 7231 (Star Formation in Orion) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of BN-IRC2 which is located in the Orion Nebula. Images of H II, Fe II, and Paschen_alpha observations will be made to characterize processes associated with star formation in the Orion Nebula. The observations were completed as scheduled, and no anomalies were reported. 1.9 Completed STIS/CCD/NIC/1/2/3 7234 (Fe II AND Mg II Emission From Redshift 4 QSOs) The Space Telescope Imaging Spectrograph (CCD) and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the redshift 4 QSO object Q0000-26. This QSO will be observed with both NICMOS grisms A and B to measure the Fe II bump and both the Fe II and Mg II 2800 A0 features, as well as the 1909 A0 C III line. The sensitivity is adequate to measure the broad Fe II bump in one orbit per source. While operating this proposal, the NICMOS suspend described in 2.3 occurred. No further anomalies were reported. 1.10 Completed Three Sets of NIC/1/2/3 7217 (GRISM Survey of Three Young Clusters) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the young clusters called the NGC-2264 and NGC-2024. Although photometric surveys have been used to great success in identifying members of embedded star clusters, significant uncertainty surrounds the mapping of broadband photometry with masses and ages of the stars. As described in 2.1, the acquisition and re-acquisition for the first iteration of this proposal failed, and the take data flag remained down for 42 observations. Ot herwise, the remaining observations were completed as planned, and further no problems were reported. 1.11 Completed NIC/1/2/3 7368 (Infrared H_2 imaging of the highly symmetric proto-stellar jets HH212 and HH111) The Near Infrared Camera (Cameras One, Two and Three) was used to make infrared observations of the jets on the Herbig-Harow Object HH212-Cent to determine the structure of faint reflection nebulae seen around the jet driving sources. This is to determine whether these are cavities excavated by the molecular outflows, or the upper surface of a large-scale disk around the source. The observations were complete as planned, and no anomalies were reported. 1.12 Completed NIC/2 7228 (Imaging and Polarimetry of YSO'S) The Near Infrared Camera (Camera Two) was used to investigate the young stellar object HH1-2-1 with coronagraphic imaging, polarimetry, and line imaging. The observations were completed as planned, and no problems were reported. 1.13 Completed NIC/1/2/3 7224 (Photometry of the 3C324 Cluster and Line/Off Line of the Radio galaxy) The Near Infrared Camera (Cameras One, Two and Three) was used to make photometric observations of the radio galaxy cluster 3C324P2. Spatially-resolved IR photometry of this aligned radio galaxy and its surrounding cluster of E class galaxies should be an important step in our understanding of the alignment effect in powerful radio galaxies and in our study of the passive evolution of the oldest systems. The observations were completed as scheduled, and no problems were reported. 1.14 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/CCD 7695 (NICMOS Grism Wavelength Calibration) The Space Telescope Imaging Spectrograph (CCD) was used to perform a grism wavelength calibration by making observations of standard object HB12. A dark calibration image was also taken. The observations were completed as scheduled, and no problems were reported. 1.16 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.17 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-4373, NGC-404, NGC-6946 and NGC-221 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.19 Completed Three Sets of NIC/1/2/3 7235 (Observations of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Hubble Deep Field area HDF-123645. This program observes a portion of the Hubble Deep Field with all three NICMOS cameras in the F160W and F110W filters to produce infrared images equivalent in depth to the existing WF/PC-2 HDF images. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed Three Sets of STIS/CCD 7918 (STIS CCD Pure Parallels for NIC/3 HDF Spec) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, spectrographic observations with the prime Hubble Deep Field (HDF) observations. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as planned, and no problems were reported. 1.21 Completed STIS/CCD 7634 (CCD Flat-Field Stability and Cosmetic Defect Fraction) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD imaging flat-field stability and cosmetic defects over monthly periods though out Cycle 7. A bias calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed STIS/CCD 7900 (STIS CCD Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, observations with the NIC/3 Camera. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as scheduled, and no anomalies were reported. 1.23 Completed NIC/3 7816 (NIC/3 Photometric Zero Point) The Near Infrared Camera (Camera Three) was used to make observations of standard star CSKD12-Cal in order to study the zero point flux for a set of filters. The NIC/3 intra-pixel sensitivity will also be measured. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 38 Successful : 36 The acquisition at 016/151546Z lost fine lock twice before finally achieving it on the third try. HSTAR 6315 was written. The proposal described in 1.7 may have been affected. Per HSTAR 6318, the guide star acquisition scheduled for 017/072521Z failed when the search radius limit was exceeded. The re-acquisition scheduled at 017/085942Z also failed. Maps obtained do not verify a pointing error. The proposal detailed in 1.10 was affected. The acquisition at 019/010502Z defaulted to fine lock on FGS-3 only. HSTAR 6320 was written. The proposal described in 1.3 may have been affected. Scheduled Reacquisitions : 26 Successful : 25 2.2 FHST Updates: Scheduled : 70 Successful : 69 As described in HSTAR 6317, the update scheduled at 017/015439Z failed with no impact to science. 2.3 Operations Notes: Database 53.B was installed on the back-up rail at 017/01Z and was analyzed with the real time system throughout the midnight shift. No problems were detected and the database was then installed on the prime rail at 017/1230Z. It was then used throughout the long weekend with no problems. The STIS bad pixel table was uplinked at 017/0431Z. At 017/011006Z, NICMOS suspended while in the observe mode during the proposal described in 1.9. A status buffer message indicated that the cause was an Intel exception. The appropriate personnel were notified and a memory dump was performed. The event signature was similar to another event which occurred on day 270 of last year. NICMOS was therefore successfully recovered at 017/0701Z per Operation Request 13512. HSTAR 6316 was written. Proposals mentioned in 1.3, 1.5, and 1.6 were also affected. At 020/034023Z, the STIS EMC experienced a re-try. HSTAR 6323 was written. As described in HSTAR 6322, the STIS MAMA2 reset at 020/080052Z while the high voltage was off. There was no impact to science. Recovery will be via the SMS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH